OpenStudy (anonymous):

Evaluate the limit using L'Hospital's rule if necessary limit (1+3/x)^(x/7) x->inf

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

it isn't necessary. plug in infinity

OpenStudy (anonymous):

how did you figure that it isn't necessary?

OpenStudy (anonymous):

you use l'hopital's only if you have 0^0 , inf^inf , 0/0 or inf/inf plugging in inf gives 1^inf = 1

OpenStudy (anonymous):

i thought the l'hopital's need to be used if it's also 1^inf.. and i plugged the answer 1^inf, and the system marked it wrong

OpenStudy (anonymous):

answer is just 1

OpenStudy (anonymous):

^inf is redundant

OpenStudy (anonymous):

i also did put in 1, but it says it's wrong

OpenStudy (math&ing001):

Ln((1+3/x)^(x/7))= x/7*Ln(1+x/3) L'hopital use f(x)=Ln(1+3/x) and g(x)=7/x as functions they both ->0 You'll get lim f(x)/g(x)= lim f'(x)/g'(x)= 3/7 So lim (1+3/x)^(x/7) = exp(3/7)

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

rip that medal off my neck

OpenStudy (anonymous):

That's okay

OpenStudy (anonymous):

Thanks

OpenStudy (math&ing001):

yw